OpenStudy (anonymous):

solve for x: 9^x = e^(x+1)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[9^{x}=e ^{(x+1)}\]\[3^{2(x)}=e ^{(x+1)}\]\[2x \log(3)=(x+1)\log\]\[2xlog(3)=x \log +\log\]\[2x \log(3)-x \log =\log\]\[x(2\log(3)-1) = 1\]\[x=\frac{1}{2\log(3)-1}\]=0.835265

OpenStudy (anonymous):

check some one ansewr frist I'm not sure